OpenStudy (anonymous):

Fan and medal? Solve the system by substitution. 2.5 - 3y = -13 3.25 - y = -14

OpenStudy (anonymous):

okay so what all do you understand about substitution?

OpenStudy (anonymous):

Nothing whatsoever ._. I suck at math and I'm really behind from my class.

OpenStudy (anonymous):

alright substitution is easy first you need to multiply the bottom by negative 3

OpenStudy (anonymous):

then you need to add the equations together

OpenStudy (anonymous):

so it will be 2.5x - 3y = 13 +9.75x +3y = 42

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

what do you get when you add them together?

OpenStudy (anonymous):

12.25x + 6y + 55

OpenStudy (anonymous):

=*

OpenStudy (anonymous):

does it say anywhere what y or x equals?

OpenStudy (anonymous):

No, I posted everything

OpenStudy (anonymous):

wait i accidentally added X's to it it should be 12.25 + 6y = 55

OpenStudy (anonymous):

Oh, okay. So don't add the x's?

OpenStudy (anonymous):

well i looked at the equation and saw that there weren't X's but add them either way

OpenStudy (anonymous):

Oh. Okay, gotcha. So what next?

OpenStudy (anonymous):

so you can subtract 12.25 from the left and from the right leaving you with 6x=42.75

OpenStudy (anonymous):

and you can divide the 42.75 by 6 leaving you with x=12.25

OpenStudy (anonymous):

X should be Y sorry

OpenStudy (anonymous):

anyways your answer should be y = 12.25

OpenStudy (anonymous):

Oh, woe. Okay! Thank you :)

OpenStudy (anonymous):

no problem :D